OpenStudy (ashleyy_smith):

A total of 430 tickets were sold for the school play. They were either adult tickets or student tickets. There were 70 fewer student tickets sold than adult tickets. How many adult tickets were sold?

rishavraj (rishavraj):

the pattern is same as tht of ur previous question :)) think over it ;)

OpenStudy (ashleyy_smith):

awwwww :( @rishavraj

OpenStudy (ashleyy_smith):

heeeellppp meee ;-;

rishavraj (rishavraj):

yup :) and sorry going to college :P

OpenStudy (ashleyy_smith):

;-;

rishavraj (rishavraj):

@Agl202 @rebeccaxhawaii I gotta go ..... so if u can

rishavraj (rishavraj):

see I am done.... so let the no of adult tickets sold be = x and u see 70 tickets fewer were sold (student tickets) so no of student tickets sold = x - 70 and also total no of tickets sold = 430

OpenStudy (ashleyy_smith):

so I subtract 430 - 70

OpenStudy (ashleyy_smith):

@mathmath333 Can u help me pls?

OpenStudy (mathmath333):

let the adult tickeys be a, let the students ticket be s given a+s=430------1.) and s=a-70 so s-a=70-------2.) solve equation 1.) and 2.)

OpenStudy (ashleyy_smith):

Do I subtract 430 and 70?

OpenStudy (ashleyy_smith):

@mathmath333

OpenStudy (mathmath333):

i think u havent studied addinng and substracting variables

OpenStudy (mathmath333):

a+s+s-a=430+70 2s=500

OpenStudy (mathmath333):

s=250 s-a=70 250-a=70 a=250-70=180 \(\Large a=180\)

OpenStudy (ashleyy_smith):

Thank u ;-;